OpenStudy (anonymous):

Find parametric equations for the rectangular equation x2 + y2 – 25 = 0.

terenzreignz (terenzreignz):

Yike... this isn't really straightforward... First... x² + y² = 25

Directrix (directrix):

For conversion to polar purposes, x = r cos theta y = r sin theta x² + y² = 25 (r cos theta)^2 + (r sin theta)^2 = 25 r² * cos² (theta) + r² * sin² (theta) = 25 r² [(cos² (theta) + sin² (theta)] = 25 r² * 1 = 25 r² = 25 r = 5 or r = -5 Note: cos² (theta) + sin² (theta) = 1 (trig identity)